Circular No. 6485

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

MICROLENSING EVENT IN LARGE MAGELLANIC CLOUD The MACHO collaboration (cf. IAUC 6312, plus D. Minniti) reports

the discovery of a likely gravitational microlensing event in progress toward the Large Magellanic Cloud: “The star in question is located at α = 5^{h}34^{m}44^{s}.4, δ = −70^{o}25′07′′ (equinox J2000.0). The star was constant at R = 19.6, V = 19.6 during 1993–1995 and has brightened by 0.6 mag during 1996 Sept. The color appears unchanged. The current best microlensing fit to Mt. Stromlo data predicts that the star should reach a peak of R and V < 18.8 between Oct. 9 and Nov. 28, and then decline symmetrically. This fit indicates an Einstein-diameter crossing time of 120 ± 30 days, which would contribute a significant optical depth to microlensing towards the LMC. Accurate photometry and low-resolution spectroscopy of this star at regular intervals over the next few weeks will be valuable to test the microlensing interpretation. Photometric measurements may possibly detect light-curve deviations that are due to a breakdown of the usual point-source, point-lens, or constant-velocity assumptions. If these exotic effects are detected, it is usually possible to estimate the distance to the lensing object. A finding chart is available on the MACHO Alert web page (cf. IAUC 6312, 6361) or ftp site darkstar.astro.washington.edu (directory macho/Alert/96-LMC-2). Potential observers are requested to contact macho@astro.washington.edu, or A. Becker at telephone 206-543-1979 or D. Bennett at telephone 219-631-8298, to coordinate observations.”

COMET C/1996 Q1 (TABUR) M. Womack and D. Suswal, Pennsylvania State University at Erie,

report on observations of HCN and CO in comet C/1996 Q1 with the National Radio Astronomy Observatory’s 12-m telescope at Kitt Peak: “We have detected the HCN J=1-0 rotational line at 89 GHz on Oct. 6 and 7 UT. The integrated line-flux area is 0.068 ± 0.009 K km/s (antenna temperature scale). The FWHM linewidth is ∼ 1.5 km/s and the line center is not measurably shifted from the ephemeris velocity. Assuming a rotational temperature of 50 K, we derive an HCN-production rate of Q(HCN) ≃ 2.5 × 10^{25} mol/s at r = 1 AU. The CO J=1-0 line could not be detected down to a 3σ level of 0.04 K km/s.”

Naked-eye m_{1} estimates: Sept. 27.12 UT, 5.7 (M. V. Zanotta, Mount Generoso, Italy); Oct. 6.20, 5.3 (A. Pereira, Cabo da Roca, Portugal); 7.20, 5.2 (C. Vitorino, Cabo da Roca, Portugal).

1996 October 8  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green